The work done in turning a magnet of magnetic moment M by an angle of 90° from the meridian, is n times the corresponding work done to turn it through an angle of 60°. The value of n is given by    


(a)    2    


(b)    1


(c)    0.5   


(d)    0.25


Grade:11

1 Answers

Aditi Chauhan
10 years ago

(a)

Magnetic moment = M; Initial angle through which magnet is turned (q1) = 90° and final angle through which magnet is turned (q2) = 60°. Work done in turning the magnet, through 90o(W1) = MB (cos 0° - cos 90°)

= MB (1-0) = MB.

Similarly, W2 = MB (cos 0° - cos 60°)